Qn 6. Let A, B, C and D be four points on a circle, taken in such a way that

the segments AC and BD have an intersection E. If AE = πDE, compute,

providing your working, the ratio between the areas of the triangles △AEB

and △CED, that is,

A(△AEB)

A(△CED)


1
Expert's answer
2022-02-28T17:03:51-0500

"\angle A= \angle D" ( angle subtended by an arc BC)


"\angle DEC= \angle AEB" ( Vertically opposite angles)


"\therefore" Triangle AED is similar to triangle DEC


Linear scale factor "=\frac {AE}{DE}=\pi"


"\implies" Area scale factor"= \pi^2"

Area of triangle AEB to area of triangle CED


"= \pi^2:1"